# UniREditBench: A Unified Reasoning-based Image Editing Benchmark

### v1

@@ 1 Introduction
Benchmark Size Reference Images Real World Scenario Game World Scenario Attribute Temporal Pose Spatial Motion Mechanic Medium Logical Long-planing Strategic Spatial (S-Obj) (S-Obj) (S-Obj) (M-Obj) (M-Obj) (M-Obj) (M-Obj) SmartEdit [11] 219 219 RISE [52] 360 70 KRIS [41] 1,267 50 UniREditBench 2,700 2,700
Recent advances in multimodal generative models have led to remarkable improvements in instruction-conditioned image editing. Generative models [40, 47, 50, 3, 33, 38, 44], including Step1X-Edit [20], FLUX-Kontext [2], Bagel [6], Nano Banana [7], and GPT-4o [12], have demonstrated a powerful ability to understand diverse textual instructions and generate semantically consistent image edits. In parallel, reinforcement learning-based training strategies [43, 18, 36, 31] are continuously advancing, further enhancing the capabilities of image editing models. With these rapid developments, the need for a more comprehensive benchmark to evaluate model editing capabilities across different aspects has become increasingly essential. Early benchmarks [20, 46] focus on local details or global stylistic changes, e.g., style transfer, color alteration, and object removal. However, they fail to cover editing tasks that require models to perform implicit reasoning [49, 8], which are commonly used in real-life applications. As illustrated in Fig. 1, when editing instructions involving real-world or human-defined game rules, current models often generate results that lack physical plausibility. To this end, recent efforts have introduced reasoning-aware evaluation across temporal, spatial, and logical dimensions [52], and proposed a knowledge-grounded taxonomy assessing factual, conceptual, and procedural knowledge types [41].
Despite their effectiveness, these benchmarks still face two significant challenges: (1) they primarily focus on single-object attribute changes in realistic scenarios, neglecting multi-object interactions and game-world scenarios involving human-defined rules (see Tab. 1). This narrow scope restricts their ability to evaluate how effectively models generalize across a wider range of complex reasoning contexts; Additionally, (2) they mainly rely on textual reference to evaluate the generated images [52, 41], which may lead to systematic misjudgments, especially in complex reasoning-based editing scenarios (see Fig. 2).
@@
To this end, this work proposes UniREditBench, a unified benchmark for reasoning-based image editing assessment with broader evaluation dimension coverage and robust evaluation pipeline. Specifically, (1) we adopt a scenario-to-category hierarchical dimension design, covering diverse reasoning types in both real-world and game-world scenarios (shown in Fig. 1): it includes 2,700 carefully curated samples organized across 8 primary dimensions and 18 sub-categories, e.g., multi-object interaction in real world, and long-horizon game planning in game world. Meanwhile, (2) as illustrated in Fig. 2, in contrast to existing work that relies solely on textual references for evaluation, we introduce additional reference GT images to facilitate direct visual comparison with the generated image. By utilizing the visual cues provided by the reference image, the evaluator is able to able to more accurately and reliably assess the alignment of the generated image with the given instruction, as shown in Fig. 2 (b). Furthermore, to ensure the diversity and reliability of samples in this benchmark, we design a multi-scenario data synthesis pipeline. Specifically, as shown in Fig. 3, (a) For real-world scenarios, we first handcraft a few reference text prompts, including the original image description, the editing instruction, and the textual reference of edited effect. These prompts are then scaled up using the VLM. Finally, all resulted textual descriptions are directly used to generate pairs of original and edited image. (b) For game-world scenarios, we first design diverse game problems, and then use Python programs to generate image pairs, instructions, and textual reference of edited effects, ensuring both logical and visual correctness in these rule-intensive scenarios [16, 27]. Ultimately, all data samples in UniREditBench undergo VLM-based filtering and human inspection to ensure their reliability and accuracy.
Based on our data synthesis pipeline, we also propose UniREdit-Data-100K, a comprehensive reasoning-based image editing dataset with high-quality chain-of-thought (CoT) reasoning annotations, consisting of detailed, step-by-step reasoning traces generated using VLM, as shown in Fig. 3. To validate its reliability and effectiveness, we fine-tune the Bagel [6] on this dataset, resulting in UniREdit-Bagel. Experimental results demonstrate that the fine-tuned model achieves substantial improvements on both UniREditBench and other out-of-distribution benchmarks [52, 41]. Additionally, through comprehensive evaluation of both open- and closed-source editing models on our UniREditBench, we reveal their strengths and weaknesses across diverse reasoning-based scenarios.
@@ 2 Related Work
Instruction-based Image Editing. Instruction-based image editing models aim to bridge semantic understanding of instructions with accurate visual manipulation. Traditional methods perform editing by altering the diffusion trajectory without requiring additional training, including partial denoising from intermediate SDE steps [21], cross-attention control [10, 30], mask-guided blending [38, 1, 33], CLIP- or diffusion-guided manipulation [14], and latent inversion for fidelity preservation [32, 13]. Besides, several studies employ visual-language models (VLMs) to provide prompts, spatial priors, or synthetic supervision to guide a generative editing model [3, 50, 51, 9]. Recent unified frameworks aim to use a single model for both image understanding and editing in a complementary direction [44, 40, 29]. For instance, Bagel [6] features a think mode that produces reasoning text prior to editing to enhance instruction fidelity and consistency. While effective, current methods still face challenges with complex reasoning-based editing, underscoring the need for comprehensive benchmarks to assess their performance across various reasoning scenarios.
@@
Reasoning-based Benchmarks for Image Generation and Editing. In T2I generation, several benchmarks [34, 25, 17, 22] have been developed to assess the reasoning capabilities of models in generating images. For example, WISE [22] focuses on assessing models’ world knowledge, such as cultural and physical understanding, while UniGenBench++ [34] unifies semantic generation evaluation, covering 10 primary dimensions and 27 sub-dimensions, such as logic reasoning, relational understanding, supporting multilingual and varying-length assessments. In image editing evaluation, recent reasoning-based benchmarks like RISEBench [52] aim to examine temporal, spatial, and logical editing capabilities of editing models. Besides, KRISBench [41] introduces a knowledge-grounded taxonomy covering factual, conceptual, and procedural types. However, these benchmarks primarily focus on single-object knowledge and attribute reasoning. We suppose that extending evaluation to multi-object interactions and scenarios governed by human-defined rules is a crucial next step. As for image quality evaluation [45, 15], recent works like UnifiedReward [37, 35] adopt the “VLM-as-a-judge” paradigm, leveraging the powerful capabilities of VLMs to score and provide explanatory judgments. In image editing tasks, evaluation is more challenging because the evaluator needs to assess not only image quality but also understand complex editing instructions and final edited effects. Most studies like RISEBench and KRISBench utilize the property model [12], to rate instruction following, temporal consistency, and image quality. Despite effectiveness, their evaluation relies solely on textual references, which may lead to systematic misjudgments in complex reasoning tasks.
@@ 3.3 Dual-Reference Evaluation
  • Instruction Following measures how accurately the generated image reflects the input instruction, focusing on whether the explicit effect of the edits is properly manifested. Here, the VLM compares the output image $G$ against both the textual reference of edited effect $R_{t}$ and the corresponding reference GT image $R_{i}$ to verify compliance:
@@
  • Visual Quality evaluates the realism and perceptual integrity of the generated output, checking for artifacts, distortions, and physical or logical implausibility in the final image:
@@
    We choose GPT-4.1 [12] as VLM evaluator. Each score $S$ ranges from 1 to 5, following prior detailed scoring guidelines [52, 41]. Finally, the overall evaluation score aggregates these via weighted sum:
@@
FLUX-Kontext-Pro 45.68 45.65 42.85 40.17 43.59 43.54 44.03 49.80 40.07 44.36 43.72
Seedream4.0 69.54 73.13 67.88 62.40 68.24 39.27 43.54 43.79 51.91 44.63 57.05
Wan2.5 74.13 69.92 64.23 65.16 68.36 63.39 46.13 56.32 52.24 54.52 60.59
Nano Banana 77.54 78.45 72.41 77.39 76.45 66.26 56.83 56.35 65.85 61.32 68.38
GPT-4o 83.27 83.93 80.43 77.62 81.31 78.82 50.28 66.02 65.66 65.19 71.64
@@
MagicBrush 43.10 46.24 43.88 47.70 45.23 63.43 33.77 30.83 35.17 40.80 40.98
Omnigen2 54.18 57.75 52.87 53.52 54.58 70.78 27.61 37.11 24.69 40.05 43.96
Step1X-Edit 59.50 57.37 56.07 56.75 57.42 62.90 34.17 44.63 44.03 46.43 50.12
Bagel-Think 61.27 59.83 53.95 57.11 58.04 65.65 43.46 43.82 40.18 48.28 51.25
Qwen-Image-Edit 75.19 73.16 71.80 67.54 71.92 57.03 34.80 47.37 37.70 44.22 56.46
UniREdit-Bagel (Ours) 79.24 78.61 76.43 71.81 76.52 84.93 73.98 86.14 84.01 82.27 78.87
@@ 3.4 Multi-Scenario Data Synthesis
For Real-World Scenario, we employ a “text-then-image” data generation strategy. Specifically, (1) this process begins with hand-crafted textual triples that describe the original image, the editing instruction, and the textual reference of edited effect (a reasoning-based narrative of the anticipated outcome). Next, we use the powerful VLM [5], to expand this initial set into a large corpus of text triples. Subsequently, (2) these curated textual triples are input to GPT-4o [12] to synthesize the original and edited images in alignment with the described textual reference of edited effect. (3) Finally, in the quality filtering stage, VLM [5] is used to assess the generated images based on visual fidelity, instruction alignment, and potential hallucination risks. Additionally, it generates reasoning chain-of-thought (CoT) text for each qualified instance, ensuring the production of high-quality, reasoning-based image editing training data.
In Game-World Scenario, game states are inherently well-suited to be represented as structured reasoning-based editing data, where instructions can naturally be solved using Python code. Inspired by Game-RL [48], (1) we first design a diverse collection of game-based problems and develop corresponding Python programs tailored to each category. (2) Then, these programs automatically generate paired original and edited images, along with instructions, textual reference effects, and programmatic CoT reasoning traces. (3) To bridge the gap between programmatic and natural language CoT reasoning formats, we use VLMs to convert these reasoning traces into explanations that align with human inference patterns. Finally, quality filtering are applied to ensure the integrity and reliability of the data.
Overall, this multi-scenario data synthesis pipeline generates our UniREditBench, a unified reasoning-based image editing benchmark, and UniREdit-Data-100K, a large-scale synthetic dataset with high-quality CoT annotations. We will provide a detailed elaboration of this dataset in the following section.
@@ 4.1 Statistical Analysis
UniREdit-Data-100K is designed with an emphasis on balance and diversity, ensuring that each reasoning category contains over 4,000 instances to effectively support model training across a wide range of editing tasks. It is divided into two primary scenario types: (i) Real-World Scenario, which captures natural object attributes and complex multi-object interactions, and (ii) Game-World Scenario, presenting structured, rule-based editing challenges, such as puzzles and strategic planning games. We visualize the word cloud for both real-world and game-world subsets in Fig. 5 (a) and the detailed distribution of samples across different categories in Fig. 5 (b). These visualizations highlight the extensive vocabulary of our dataset that captures the diverse visual attributes, as well as its broad coverage across various categories.
@@ 4.2 UniREdit-Bagel
To further validate the effectiveness of our dataset, we use it to fine-tune Bagel [6], a unified understanding and generative model. Specifically, each training sample consists of the input image $O$, an editing instruction $I$, a stepwise CoT text $C$ that grounds the edit effects step by step, and the target edited image $G$. During training, the original image and instruction are first input into the model, which then generates a textual reasoning trace and synthesizes the edited image. We supervise both the textual reasoning trace and the visual edit. Formally, for reasoning text supervision, we minimize the negative log-likelihood:
@@
For image generation, we supervise the latent flow-matching loss [19] between the VAE latents of $O$ and $G$, conditioned on $(O,I,C)$:

### v2

@@ 1 Introduction
Benchmark Size Reference Images Real World Scenario Game World Scenario Attribute Temporal Pose Spatial Motion Mechanic Medium Logical Long-planing Strategic Spatial (S-Obj) (S-Obj) (S-Obj) (M-Obj) (M-Obj) (M-Obj) (M-Obj) SmartEdit [huang2024smartedit] 219 219 RISE [zhao2025envisioning] 360 70 KRIS [wu2025kris] 1,267 50 UniREditBench 2,700 2,700
Recent advances in multimodal generative models have led to remarkable improvements in instruction-conditioned image editing. Generative models [wu2025omnigen2, yu2025anyedit, zhang2023magicbrush, brooks2023instructpix2pix, wang2024primecomposer, wang2025dreamtext, xin2025lumina], including Step1X-Edit [liu2025step1x], FLUX-Kontext [flux_kontext], Bagel [deng2025bagel], Nano Banana [Gemini-2.5-Flash-Image], and GPT-4o [hurst2024gpt], have demonstrated a powerful ability to understand diverse textual instructions and generate semantically consistent image edits. In parallel, reinforcement learning-based training strategies [xiao2025mindomni, li2025uniworld, wang2025pref, wang2024lift] are continuously advancing, further enhancing the capabilities of image editing models. With these rapid developments, the need for a more comprehensive benchmark to evaluate model editing capabilities across different aspects has become increasingly essential. Early benchmarks [liu2025step1x, ye2025imgedit] focus on local details or global stylistic changes, e.g., style transfer, color alteration, and object removal. However, they fail to cover editing tasks that require models to perform implicit reasoning [zhang2025r, he2025reasoning], which are commonly used in real-life applications. As illustrated in Fig. 1, when editing instructions involving real-world or human-defined game rules, current models often generate results that lack physical plausibility. To this end, recent efforts have introduced reasoning-aware evaluation across temporal, spatial, and logical dimensions [zhao2025envisioning], and proposed a knowledge-grounded taxonomy assessing factual, conceptual, and procedural knowledge types [wu2025kris].
Despite their effectiveness, these benchmarks still face two significant challenges: (1) they primarily focus on single-object attribute changes in realistic scenarios, neglecting multi-object interactions and game-world scenarios involving human-defined rules (see Tab. 1). This narrow scope restricts their ability to evaluate how effectively models generalize across a wider range of complex reasoning contexts; Additionally, (2) they mainly rely on textual reference to evaluate the generated images [zhao2025envisioning, wu2025kris], which may lead to systematic misjudgments, especially in complex reasoning-based editing scenarios (see Fig. 2).
@@
To this end, this work proposes UniREditBench, a unified benchmark for reasoning-based image editing assessment with broader evaluation dimension coverage and robust evaluation pipeline. Specifically, (1) we adopt a scenario-to-category hierarchical dimension design, covering diverse reasoning types in both real-world and game-world scenarios (shown in Fig. 1): it includes 2,700 carefully curated samples organized across 8 primary dimensions and 18 sub-categories, e.g., multi-object interaction in real world, and long-horizon game planning in game world. Meanwhile, (2) as illustrated in Fig. 2, in contrast to existing work that relies solely on textual references for evaluation, we introduce additional reference GT images to facilitate direct visual comparison with the generated image. By utilizing the visual cues provided by the reference image, the evaluator is able to more accurately and reliably assess the alignment of the generated image with the given instruction, as shown in Fig. 2 (b). Furthermore, to ensure the diversity and reliability of samples in this benchmark, we design a multi-scenario data synthesis pipeline. Specifically, as shown in Fig. 3, (a) For real-world scenarios, we first handcraft a few reference text prompts, including the original image description, the editing instruction, and the textual reference of edited effect. These prompts are then scaled up using the VLM. Finally, all resulted textual descriptions are directly used to generate pairs of original and edited image. (b) For game-world scenarios, we first design diverse game problems, and then use Python programs to generate image pairs, instructions, and textual reference of edited effects, ensuring both logical and visual correctness in these rule-intensive scenarios [li2024vcbench, tong2025code2logic]. Ultimately, all data samples in UniREditBench undergo VLM-based filtering and human inspection to ensure their reliability and accuracy.
Based on our data synthesis pipeline, we also propose UniREdit-Data-100K, a comprehensive reasoning-based image editing dataset with high-quality chain-of-thought (CoT) reasoning annotations, consisting of detailed, step-by-step reasoning traces generated using VLM, as shown in Fig. 3. To validate its reliability and effectiveness, we fine-tune the Bagel [deng2025bagel] on this dataset, resulting in UniREdit-Bagel. Experimental results demonstrate that the fine-tuned model achieves substantial improvements on both UniREditBench and other out-of-distribution benchmarks [zhao2025envisioning, wu2025kris]. Additionally, through comprehensive evaluation of both open- and closed-source editing models on our UniREditBench, we reveal their strengths and weaknesses across diverse reasoning-based scenarios.
@@ 2 Related Work
Instruction-based Image Editing. Instruction-based image editing models aim to bridge semantic understanding of instructions with accurate visual manipulation. Traditional methods perform editing by altering the diffusion trajectory without requiring additional training, including partial denoising from intermediate SDE steps [meng2021sdedit], cross-attention control [hertz2022prompt, wang2023enhancing], mask-guided blending [wang2025dreamtext, avrahami2022blended, wang2024primecomposer], CLIP- or diffusion-guided manipulation [kim2022diffusionclip], and latent inversion for fidelity preservation [wang2024magicface, kawar2023imagic]. Besides, several studies employ visual-language models (VLMs) to provide prompts, spatial priors, or synthetic supervision to guide a generative editing model [brooks2023instructpix2pix, zhang2023magicbrush, zhang2024hive, he2024freeedit, han2025controlthinker]. Recent unified frameworks aim to use a single model for both image understanding and editing in a complementary direction [xin2025lumina, wu2025omnigen2, wang2024emu3]. For instance, Bagel [deng2025bagel] features a think mode that produces reasoning text prior to editing to enhance instruction fidelity and consistency. While effective, current methods still face challenges with complex reasoning-based editing, underscoring the need for comprehensive benchmarks to assess their performance across various reasoning scenarios.
@@
Reasoning-based Benchmarks for Image Generation and Editing. In T2I generation, several benchmarks [wang2025unigenbench++, sun2025t2i, li2025easier, niu2025wise] have been developed to assess the reasoning capabilities of models in generating images. For example, WISE [niu2025wise] focuses on assessing models’ world knowledge, such as cultural and physical understanding, while UniGenBench++ [wang2025unigenbench++] unifies semantic generation evaluation, covering 10 primary dimensions and 27 sub-dimensions, such as logic reasoning, relational understanding, supporting multilingual and varying-length assessments. In image editing evaluation, recent reasoning-based benchmarks like RISEBench [zhao2025envisioning] aim to examine temporal, spatial, and logical editing capabilities of editing models. Besides, KRISBench [wu2025kris] introduces a knowledge-grounded taxonomy covering factual, conceptual, and procedural types. However, these benchmarks primarily focus on single-object knowledge and attribute reasoning. We suppose that extending evaluation to multi-object interactions and scenarios governed by human-defined rules is a crucial next step. As for image quality evaluation [xu2023imagereward, kirstain2023pick], recent works like UnifiedReward [wang2025unified, wang2025unified-think] adopt the “VLM-as-a-judge” paradigm, leveraging the powerful capabilities of VLMs to score and provide explanatory judgments. In image editing tasks, evaluation is more challenging because the evaluator needs to assess not only image quality but also understand complex editing instructions and final edited effects. Most studies like RISEBench and KRISBench utilize the property model [hurst2024gpt], to rate instruction following, temporal consistency, and image quality. Despite effectiveness, their evaluation relies solely on textual references, which may lead to systematic misjudgments in complex reasoning tasks.
@@ 3.3 Dual-Reference Evaluation
  • Instruction Following measures how accurately the generated image reflects the input instruction, focusing on whether explicit effect of the edit is properly manifested. Here, the VLM compares the output image $G$ against both the textual reference of edited effect $R_{t}$ and the corresponding reference GT image $R_{i}$ to verify compliance:
@@
  • Visual Quality evaluates the realism of the generated output, checking for artifacts, distortions, and physical or logical implausibility in the final image:
@@
    We choose GPT-4.1 [hurst2024gpt] as VLM evaluator. Each score $S$ ranges from 1 to 5, following prior detailed scoring guidelines [zhao2025envisioning, wu2025kris]. Finally, the overall evaluation score aggregates these via weighted sum:
@@
FLUX-Kontext-Pro 47.35 47.16 44.37 41.44 45.00 49.12 48.58 51.16 40.49 46.52 45.77
Seedream4.0 67.98 72.93 65.07 59.49 66.22 38.87 42.16 44.09 51.65 45.38 55.77
Wan2.5 74.66 69.74 63.51 62.87 67.23 63.73 47.13 55.00 54.93 52.67 61.36
Nano Banana 77.10 78.88 71.86 74.70 75.22 66.74 56.11 56.83 64.91 60.39 68.26
GPT-4o 82.67 84.75 79.81 77.40 81.01 77.73 51.44 67.61 63.79 62.07 73.39
@@
MagicBrush 43.97 46.09 42.93 46.64 44.69 63.58 33.59 30.72 35.31 36.85 40.77
Omnigen2 55.47 58.27 51.44 50.70 53.69 70.28 27.50 36.25 24.31 33.14 43.41
Lumina-DiMOO 52.84 52.31 50.31 50.83 51.44 61.23 36.96 39.57 53.09 45.61 48.54
Step1X-Edit 59.69 56.36 53.85 54.84 55.93 65.68 34.89 47.01 43.89 44.00 50.15
Bagel-Think 61.45 59.51 52.65 55.68 56.80 66.29 43.23 43.00 41.30 45.10 50.96
DreamOmni2 63.52 58.35 52.68 54.02 56.64 72.42 42.17 48.80 48.09 48.98 52.81
UniWorld-V2 72.96 69.37 63.65 61.69 66.55 49.27 40.00 53.41 37.53 43.19 54.87
Qwen-Image-Edit 75.68 73.03 70.59 64.67 70.95 56.73 36.63 48.80 37.68 41.92 56.52
UniREdit-Bagel (Ours) 76.73 77.80 76.57 71.44 75.74 84.90 72.83 84.88 83.72 80.48 78.15
@@ 3.4 Multi-Scenario Data Synthesis
For Real-World Scenario, we employ a “text-then-image” data generation strategy. Specifically, (1) this process begins with hand-crafted textual triples that describe the original image, the editing instruction, and the textual reference of edited effect (a reasoning-based narrative of the anticipated outcome). Next, we use the powerful VLM [comanici2025gemini], to expand this initial set into a large corpus of text triples. Subsequently, (2) these curated textual triples are input to GPT-4o [hurst2024gpt] to synthesize the original and edited images in alignment with the described textual reference of edited effect. (3) Finally, in the quality filtering stage, VLM [comanici2025gemini] is used to assess the generated images based on visual fidelity, instruction alignment, and potential hallucination risks. Additionally, it generates reasoning chain-of-thought (CoT) text for each qualified instance, ensuring the production of high-quality, reasoning-based image editing training data.
In Game-World Scenario, game states are inherently well-suited to be represented as structured reasoning-based editing data, where instructions can naturally be solved using Python code. Inspired by Game-RL [tong2025code2logic], (1) we first design a diverse collection of game-based problems and develop corresponding Python programs tailored to each category. (2) Then, these programs automatically generate paired original and edited images, along with instructions, textual reference effects, and programmatic CoT reasoning traces. (3) To bridge the gap between programmatic and natural language CoT reasoning formats, we use VLMs to convert these reasoning traces into explanations that align with human inference patterns. Finally, quality filtering are applied to ensure the integrity and reliability of the data.
Overall, this multi-scenario data synthesis pipeline generates UniREditBench, a unified reasoning-based image editing benchmark, and UniREdit-Data-100K, a large-scale dataset with high-quality CoT annotations. We detail the elaboration of this dataset in the next section.
@@ 4.1 Statistical Analysis
UniREdit-Data-100K is designed with an emphasis on balance and diversity, ensuring that each reasoning category contains over 4,000 instances to effectively support model training across a wide range of editing tasks. It is divided into two primary scenarios: (i) Real-World Scenario, which captures natural object attributes and complex multi-object interactions, and (ii) Game-World Scenario, presenting structured, rule-based editing challenges, such as puzzles and strategic planning games. We visualize the word cloud for both real-world and game-world subsets in Fig. 5 (a) and the detailed distribution of samples across different categories in Fig. 5 (b). These visualizations highlight the extensive vocabulary as well as the broad coverage across various categories of our dataset.
@@ 4.2 UniREdit-Bagel
To further validate the effectiveness of our dataset, we use it to fine-tune Bagel [deng2025bagel], a unified understanding and generative model. Specifically, each training sample consists of the input image $O$, an editing instruction $I$, a stepwise CoT text $C$ that grounds the edit effects step by step, and the target edited image $G$. During training, the original image and instruction are first input into the model, which then generates a textual reasoning trace and synthesizes the edited image. We supervise both the textual reasoning trace and the visual edit. Formally, for reasoning text supervision, we minimize the negative log-likelihood:
@@
For image generation, we supervise the latent flow-matching loss [lipman2022flow] between the VAE latents of $O$ and $G$, conditioned on $(O,I,C)$: